OpenStudy (anonymous):

Find the domain of f(x) = ln(5 − x).

OpenStudy (anonymous):

x<5 and x<0

OpenStudy (anonymous):

From the definition of ln \[e^y=5-x\] (if you subbed any number into y for this equation you would never get zero or a negative) From this definition you can see that 5-(x) > 0 Therefore you can deduce that any positive number greater than 5 that you place into this equation you will end up with a negative. Although if you insert anything close to anything below or close to 5 you will end up with a positive. for example -1 5-(-1) > 0 a negative multiplied by a negative is a positive thus you have 6 5+1 = 6 and 6>0 Thus your domain of definition is (negative infinity, 5) please note that I used round brackets instead of square brackets because the number set does not include 5 as the equation cannot be equal to zero. With infinities always use round brackets. Hope this is easy enough for you